True or False: The freezing point of a 0.15 MgCl2 (aq) should theoretically be lower than the freezing point of a 0.25m KI (Aq). Explain why
1
Expert's answer
2019-09-17T04:57:42-0400

No. It is not true.

The lowering of a freezing point is expressed as:


"\delta T=iC_mK"

,where i - vant Hoff factor, which is equal to the number of ions produced by one formula unit of the salt (if the dissociation is complete). Thus, i(MgCl2) = 3 and i(KI) = 2:


"MgCl_2 \rightarrow Mg^{+2}+2Cl^-"

"KI \rightarrow K^++I^-"


Finally,


"\delta T(MgCl_2)=3*0.15*K=0.45K"

"\delta T(KI)=2*0.25*K=0.50K"